OpenStudy (anonymous):

Help please! Jack and Brenda Bushnell pay $2,870.19 in annual property taxes. The rate of assessment for their tax district is 54%, and their tax rate is 39.87 mills. What is their property’s market value to the nearest cent?

OpenStudy (anonymous):

I don't know what I am doing wrong, but this is what I get and apparently I'm incorrect. 2870.19= .03987*(m*.54) 2870.19=0.0215(m) 2870.19/0.0215=m 133497.209=m 133497.21=m